OpenStudy (anonymous):

v^2 = 2GM / d, solve for M. So far I have .5v^2/G = M/d but stuck from there.

OpenStudy (agent0smith):

If you've gotten it to \[\Large \frac{ 0.5 v^2 }{ G } = \frac{ M }{d }\] then the last step is easy... the M is being divided by d, so what do you need to do to both sides of the equation to get M?

OpenStudy (anonymous):

multiply by d?

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

should have been obvious! lol. Thank you!

OpenStudy (agent0smith):

haha, it happens. I was surprised you got to where you did, and then got stuck!